# SpargeAttn: Accurate Sparse Attention Accelerating Any Model Inference

### v2

@@ 4.2 Quality and Efficiency Evaluation
End-to-end metrics. We assess the end-to-end metrics of various models using SpargeAttn compared to using full attention and baselines. Table 1 shows the results. We can observe that our method incurs almost no end-to-end metric loss across various models compared to Full-Attention and surpasses baselines with various sparsity levels in terms of end-to-end accuracy. Fig. 6 and 7 show some visible comparison examples on Flux, Stable-Diffusion3.5, and Mochi, showing that SpargeAttn incurs no performance loss and outperforms baselines.
Figure 9: Kernel speed comparison under varying sparsity on RTX4090. Input tensors have a sequence length of 22K and a head dimension of 128. SpargeAttn+FA2/Sage/Sage2 means deploying our method on FlashAttention2, SageAttention or SageAttention2 (Zhang et al., 2024a).
@@ 4.3 Ablation Study and key Insights
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$+$M_{pv}$ on Llama3.1 in NeedleInAHaystack task with 128K sequence length.
SpargeAttn enhance the LLM performance. From Table 1, Fig. 8 and  10, we observe that SpargeAttn enhances LLM performance in long-context tasks. This improvement may result from the fact that sparse attention helps the LLM focus on more relevant information.
@@ A.2 Ablation Study of Self-Similarity Judge
Figure 10: A NeedleInAHaystack comparison example on Llama3.1. The sparsity of SpargeAttn, MInference, and FlexPrefill is 0.36, 0.3, and 0.3.
Table 11: End-to-end metrics on Llama3.1 in the NeedleInAHaystack task with 16-28K sequence lengths.
@@ A.3 Additional Experiments
In this section, we present additional experimental results further to evaluate the performance of SpargeAttn compared to baselines. Fig. 10 and 11 show the results on Llama3.1 in the NeedleInAHaystack task with 16-28K sequence length. Fig 11 shows a visible comparison example on Mochi.

### v4

@@ 4.2 Quality and Efficiency Evaluation
End-to-end metrics. We assess the end-to-end metrics of various models using SpargeAttn compared to using full attention and baselines. Table 1 shows the results. We can observe that our method incurs almost no end-to-end metric loss across various models compared to Full-Attention and surpasses baselines with various sparsity levels in terms of end-to-end accuracy. Fig. 678, and 12 show some visible comparison examples on CogvideoX, Flux, Stable-Diffusion3.5, Mochi, and Open-Sora-Plan, showing that SpargeAttn incurs no performance loss and outperforms baselines.
Figure 10: Kernel speed comparison under varying sparsity. Input tensors have a sequence length of 22K and a head dimension of 128. SpargeAttn+FA2 means deploying our method on FlashAttention2.
@@ 4.3 Ablation Study and key Insights
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$+$M_{pv}$ on Llama3.1 in Needle-in-a-Haystack task with 128K sequence length.
SpargeAttn enhance the LLM performance. From Table 1, Fig. 9 and  11, we observe that SpargeAttn enhances LLM performance in long-context tasks. This improvement may result from the fact that sparse attention helps the LLM focus on more relevant information.
@@
Sparsity analysis over diffusion model. We conduct a detailed analysis of sparsity in CogvideoX across all layers, heads, timesteps, and samples using SpargeAttn to get more insights (See Appendix A.4 for detailed figures). We find that sparsity varied with layers and heads, indicating that setting different hyperparameters for each layer and head is necessary. We also find that for diffusion models, the sparsity increases with the sample timesteps.
@@ 5 Conclusion
Acknowledgment
This work was supported by the NSFC Projects (Nos. 92270001, 62376131). J.Z is also supported by the XPlorer Prize.
Impact Statement
This paper presents work that aims to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.2 Ablation Study of Self-Similarity Judge
Figure 11: A Needle-in-a-Haystack comparison example on Llama3.1. The sparsity of SpargeAttn, MInference, and FlexPrefill is 0.36, 0.3, and 0.3.
Figure 12: Visible examples on Open-sora-Plan.
Figure 13: Comparison examples on Mochi. The sparsity of SpargeAttn, MInference, and FlexPrefill is 0.47, 0.3, and 0.4.
Table 11: End-to-end metrics on Llama3.1 in the Needle-in-a-Haystack task with 16-28K sequence lengths.
@@ A.3 Additional Experiments
In this section, we present additional experimental results further to evaluate the performance of SpargeAttn compared to baselines. Fig. 11 and 11 show the results on Llama3.1 in the Needle-in-a-Haystack task with 16-28K sequence length. Fig 13 shows a visible comparison example on Mochi.
A.4 Sparsity analysis over diffusion model
In this section, we analyze the sparsity patterns in CogvideoX across different dimensions: model layers, denoising timesteps, input samples, and attention heads. Figure 14 illustrates the layer-wise sparsity. Figure 15 demonstrates timestep-wise sparsity. Figure 16 highlights sample-wise sparsity. Figure 17 presents head-wise sparsity, illustrating the diversity in attention behavior across different heads. These analyses are helpful for the design of some diffusion algorithms (Zheng et al., 2023, 2024b, 2024a, 2025; Zhao et al., 2024, 2025a; Wang et al., 2024).
Figure 14: Layer-wise sparsity of CogvideoX.
Figure 15: Timestep-wise sparsity of CogvideoX.
Figure 16: Sample-wise sparsity of CogvideoX.
Figure 17: Head-wise sparsity of CogvideoX.